OpenStudy (anonymous):

Plz Help! Can you find the interquartile range of this data set?? 76, 82, 73, 93, 91, 81, 79, 86, 84 I thought the answer was 11 but my teacher said it was wrong!

OpenStudy (amistre64):

how did you go about getting the 11? that way we can find the trouble area, if there is one

OpenStudy (amistre64):

if we subtract 70 from all the data points we get: {76, 82, 73, 93, 91, 81, 79, 86, 84}-70 6, 12, 3, 23, 21, 11, 9, 16, 14 ; which is an equivalent set

OpenStudy (anonymous):

Um I ordered the data, found the median, took out the median and then found the median of those two sets. Then I subtracted the uq from the lq.

OpenStudy (amistre64):

sounds valid, the median of the medians ... i tend to just go for broke: there are 9 data points; 9/4 = 2.x, rnd to 3 9*3/4 = 6.x, rnd to 7 on the ordered arrangment, subtract the 7th postiion from the 3rd position

OpenStudy (amistre64):

3,6,9,11,12,14,16,21,23 ^ ^ 16-9 = 7

OpenStudy (anonymous):

Oh thanks!

OpenStudy (amistre64):

most likely you just pulled off the wrong values, but the method was good